# Benchmarking Web API Integration Code Generation

### v5

@@
Abstract
API integration is a cornerstone of our digital infrastructure, enabling software systems to connect and interact. However, as shown by many studies, writing or generating correct code to invoke APIs, particularly web APIs, is challenging. Although large language models (LLMs) have become popular in software development, their effectiveness in automating the generation of web API integration code remains unexplored. In order to address this, we present WAPIIBench, a dataset and evaluation pipeline designed to assess the ability of LLMs to generate web API invocation code. Our experiments with several open-source LLMs reveal that generating API invocations poses a significant challenge, resulting in hallucinated endpoints, incorrect argument usage, and other errors. None of the evaluated open-source models was able to solve more than 40% of the tasks.
I Introduction
Web APIs provide services and functionality through a standardized, HTTP-based interface. Developers leverage them to rapidly create applications (powered by so-called API integration code) that seamlessly integrate with a wide range of online services. The market surrounding API integrations is experiencing explosive growth [1, 2].
However, writing correct API integration code is a tedious and challenging task. Developers need to understand and correctly utilize a significant amount of information to be able to write an API invocation in compliance with its specification [3, 4]. Moreover, the task of ensuring correct API usage is further complicated by the fact that APIs evolve and change over time.
Large language models (LLMs) have great potential to boost productivity in software development thanks to their ability to assist developers, e.g., by automatically generating program code from natural language descriptions [5]. However, LLMs may also hallucinate and make mistakes, raising concerns about correctness [6, 7], security [8, 9], and general quality [10] of LLM-generated code.
@@ I Introduction
While the capabilities of LLMs have been investigated for many different software engineering tasks, including library (i.e., non-web) API invocations [11], their aptitude for web API integrations appears to be rather unexplored (cf. also related work in Section IV). The question is: Given the pivotal role of web APIs, can we unlock the potential of LLMs for streamlining and accelerating the development processes of software that involves web API integration? As a first step towards answering this question, we need to understand how well existing LLMs support that task. To this end, we formulate the following research question:
RQ: How correct is LLM-generated web API invocation code, and what types of errors are commonly present in such code?
Methodology
Flow diagram visualizing the evaluation pipeline described in the caption.
Figure 1: Benchmark design for evaluating the capabilities of LLMs in generating web API invocation code. 1) Based on an API $a$ and its specifications $s$, an API invocation tasks $t$ and corresponding correct request configurations $c$ are created. 2) For each $t$, the LLM under evaluation generates an API invocation ii. 3) $i$ is executed in a controlled environment, yielding a request configuration $c^{\prime}$. 4) $c^{\prime}$ is compared to $c$ and validated against $s$ to obtain various metrics, shown in Table II.
@@
To arrive at an answer, we create WAPIIBench, a benchmark for web API integration tasks. Specifically, we perform the following steps. First, we create a dataset of web API invocation tasks and expected outcomes. Using a powerful state-of-the-art LLM, we generate the initial version of the dataset and subject it to a thorough manual review and correction process. Second, we implement an evaluation pipeline that allows us to safely execute web API invocation code and evaluate the resulting requests to obtain fine-grained correctness metrics. For this purpose, we develop a novel evaluation methodology. Third, we use our dataset and evaluation pipeline to evaluate a comprehensive set of LLMs, including models from the StarCoder [12, 13], DeepSeek-Coder [14, 15], Qwen-Coder [16], (Code) Llama [17, 18], and GPT families.
@@
Listing 1: General structure of an API invocation using Axios and JavaScript.
xxxxxx
@@
axios.<method>(’http://server.com/path/to/endpoint/arg1’, {
@@
Since web APIs can be used in many ways, we need to be more specific and focus our research specifically on the following setup: We target real-world web APIs that adhere to OpenAPI [^1] https://www.openapis.org/, the most widely used specification standard for web APIs [19]. We generate code in JavaScript, one of the most popular programming languages [20], and use Axios [^2] https://axios-http.com/, the leading JavaScript library for invoking web APIs [21], which allows explicit configuration of all components of an HTTP requests.
@@
The general structure of the API invocations we consider is shown in Listing 1. An invocation comprises an HTTP method (GET, POST, etc.), a URL with server address and (potentially parameterized) path to the endpoint, a request body, a request header, and query parameters. Such web API invocations have some unique challenges compared to function calls in a local codebase: 1. Operations are identified by the combination of a method and a long URL string, not just a simple function name; 2. there is not just one but multiple argument lists and the number of available arguments can be much higher while many arguments have complex, nested data types; and 3. web API endpoints are documented externally, limiting accessibility for LLMs, and the used specification format is much more complex than, e.g., that of documentation comments used for local functions.
Evaluation results
@@
Our study reveals that, without any in-context information about the APIs, most LLMs demonstrate some understanding of the assigned tasks, show familiarity with the general structure of invocations in JavaScript/Axios, and appear to have memorized portions of the APIs’ specifications. However, even the best open-source models generate correct invocations only 40% of the time. In particular, they frequently hallucinate endpoint URLs (up to 39%) and parameter names (up to 31%).
@@
In summary, by introducing WAPIIBench, we make the following contributions:
A first-of-its-kind dataset of web API invocation tasks
across four real-world APIs with nearly 400 pairs of tasks and expected outcomes, which can be used for execution-based correctness analysis. → Section II-A
An open-source evaluation pipeline
for automatically and safely evaluating the performance of code generation models on web API invocation tasks using fine-grained metrics. → Section II-B – Section II-D
Novel empirical insights
on the (in)ability of LLMs to generate correct API invocation code. → Section III
WAPIIBench is available on GitHub [^3] https://github.com/stg-tud/WAPIIBench. In addition, we provide all model-generated codes in our artifact [^4] https://doi.org/10.5281/zenodo.17607587. Additional details and comprehensive results are provided in an appendix at the end of this paper.
II Benchmark Design
WAPIIBench’s evaluation pipeline, depicted in Figure 1, consists of four stages: 1. dataset creation, 2. code generation, 3. code execution, and 4. correctness analysis. They are detailed in the following subsections.
II-A Dataset Creation
@@
Instead, we opted for a synthetic approach, using an LLM $\operatorname{Gen}$ and API specifications $s$ to generate task–configuration pairs for our dataset: $\operatorname{Gen}(a,s)=(t,c)$. To ensure our dataset represents real-world scenarios, we chose APIs of popular services across different domains—Asana, Google Calendar, Google Sheets, and Slack—which are specified following OpenAPI, the most commonly used industry standard [19]. As $\operatorname{Gen}$, we used Gemini 1.5 Pro [^6] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/ for its very large context window (2M tokens), given that OpenAPI specifications can be very long (e.g., Slack’s contains over 20,000 lines of text). Together with the full specification, we provided the model with very detailed (but mostly non-API-specific) instructions on how to construct $t$ and $c$ in a zero-shot prompt to avoid pitfalls of few-shot prompting [22] and increase generalizability. The prompt is provided in the appendix. In total, we synthesized 395 samples—one for each endpoint of the selected APIs. A detailed breakdown is shown in Table I.
To ensure the validity of the AI-generated data, we curated all samples with automated and manual checks. The automated checks caught any inconsistencies between the generated configurations $c$ and the specifications $s$. The 9 samples that failed these checks were manually inspected and corrected. Afterward, we manually checked all 395 samples to catch remaining issues that cannot be checked automatically, especially underspecification and ambiguities in the tasks $t$, in order to ensure that there is one unique $c$ for each $t$. Concretely, we checked 1. that $t$ is well-formulated and solvable; 2. that $c$ actually solves $t$; and 3. that every argument in $t$ is reflected in $c$ and vice versa. 58 samples had one or multiple issues and were fixed by adjusting $t$ and/or $c$.
@@
TABLE I: Number of Endpoints per HTTP Method for Each API (Each Endpoint Corresponds to One Sample in the Dataset)
API Total POST GET PUT
DELETE
PATCH
@@
II-B Code Generation
Note that the code generation task described here is considerably different from the dataset generation task from Section II-A. In the latter, Gemini had access to the full API specification and very detailed instructions, making this a task about processing in-context information. The models under evaluation, on the other hand, have to rely solely on memorized knowledge about the APIs, as we want to find out how well they can recall these details from their training data. This is a relevant concern as context length is limited, and it is often not desirable to fill the context with too much information. For instance, Slack’s specification has around 126K tokens, exceeding the context length of most models we evaluated.
Listing 2: Simplified task from our dataset with correct implementation.
xxxxxx
@@ II-B Code Generation
axios.post(
’https://www.googleapis.com/calendar/v3/calendars’, {
@@
Listing 3: Request configuration corresponding to the task in Listing 2.
xxxxxx
@@
"url":
"https://www.googleapis.com/calendar/v3/calendars",
@@
II-C Code Execution
Our approach to evaluating LLMs’ capabilities in generating API invocation code is based on functional correctness rather than syntactic similarity to a reference implementation, as the latter is prone to misjudgments [23]. To measure functional correctness, we need to execute the generated code and assess its outcome, analogous to unit testing [24].
When executing $i$, we need to address challenges unique to our setting. First, direct execution of generated code is unsafe as it could send arbitrary requests to external servers. Second, we need to capture and evaluate the configuration $c^{\prime}$ of the API request implemented by $i$, which is a side effect and not the return value of $i$. Another common challenge when evaluating code generation is excess code [25], i.e., code produced when the model terminates too late.
@@ II-C Code Execution
II-D Correctness Analysis
TABLE II: Evaluation Metrics for Correctness and Specification-Compliance of API Invocations
@@ II-D Correctness Analysis
We compare the configuration $c^{\prime}$, captured in the previous step, with the ground-truth $c$ from our dataset: $c^{\prime}\overset{?}{=}c$. However, as web API requests contain multiple components—URL, method, and parameters in different locations (cf. Listing 1)—only looking for exact matches is insufficient. Instead, we compare the configurations in an element-wise manner as well as in aggregated forms to facilitate a fine-grained evaluation. Additionally, we validate $c^{\prime}$ against the API specification $s$ to distinguish between endpoints and arguments that are undesired but specification-compliant and ones that are actually illegal. The key metrics we calculate are explained in Table II and additional ones are provided in the appendix. Note that some metrics apply only to one of the setups (full/argument completion), as mentioned in Section II-B.
III Evaluation
Using our dataset and evaluation pipeline, we study the performance of 21 state-of-the-art open-source LLMs on web API invocation code generation. We primarily evaluated models that have undergone fine-tuning on code data—CodeT5+ [26] (16B), StarCoder [12] (15.5B), StarCoder2 [13] (3B, 7B, 15B), DeepSeek-Coder [14] (1.3B, 6.7B, 7B, 33B), DeepSeek-Coder-V2 [15] (16B), Qwen2.5-Coder [16] (0.5B, 1.5B, 3B, 7B, 14B, 32B), and Code Llama [17] (7B, 13B, 70B)—and additionally Llama 3.1 [18] (8B, 70B). These models were selected due to their frequent appearance in coding leaderboards [^7] E.g., https://evalplus.github.io/leaderboard.html. To obtain an upper bound on performance, we also evaluated OpenAI’s commercial GPT-4o [^8] https://openai.com/index/hello-gpt-4o/ and GPT-4o mini [^9] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ models. For a lack of space and for more visual clarity, we show only a subset of these models in our plots and tables, but our following analysis considers all of them. Comprehensive result tables are provided in the appendix.

### v6

# Evaluating and Mitigating Errors in LLM-Generated Web API Integrations

@@
Abstract.
API integration is a cornerstone of our digital infrastructure, enabling software systems to connect and interact. However, as shown by many studies, writing or generating correct code to invoke APIs, particularly web APIs, is challenging. Although large language models (LLMs) have become popular in software development, their effectiveness in automating the generation of web API integration code remains unexplored. In order to address this, we present WAPIIBench, a dataset and evaluation pipeline designed to assess the ability of LLMs to generate web API invocation code. Our experiments with several open-source LLMs reveal that generating API invocations poses a significant challenge, resulting in hallucinated endpoints, incorrect argument usage, and other errors. None of the evaluated open-source models was able to solve more than 40% of the tasks. Motivated by those findings, we explore the potential of constrained decoding for generating API invocations. To this end, we propose an automatic translation from API specifications to constraints. Our approach prevents violations of API usage rules and significantly increases the overall correctness of the generated code, on average by 90% and 135%, depending on the provided starter code.
artificial intelligence, software engineering, large language models, code generation, web APIs, benchmarks, constrained decoding
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Automatic programmingccs: Computing methodologies Neural networksccs: Information systems RESTful web services
1. Introduction
Web APIs provide services and functionality through a standardized, HTTP-based interface. Developers leverage them to rapidly create applications (powered by so-called API integration code) that seamlessly integrate with a wide range of online services. The market surrounding API integrations is experiencing explosive growth (MarketsandMarkets, 2024; Postman, 2024).
However, writing correct API integration code is a tedious and challenging task. Developers need to understand and correctly utilize a significant amount of information to be able to write an API invocation in compliance with its specification (Robillard, 2009; Rauf et al., 2019). Moreover, the task of ensuring correct API usage is further complicated by the fact that APIs evolve and change over time.
Large language models (LLMs) have great potential to boost productivity in software development thanks to their ability to assist developers, e.g., by automatically generating program code from natural language descriptions (Cui et al., 2025). However, LLMs may also hallucinate and make mistakes, raising concerns about correctness (Dou et al., 2024; Tambon et al., 2025), security (Pearce et al., 2022; Perry et al., 2023), and general quality (GitClear, 2025) of LLM-generated code.
@@ 1. Introduction
While the capabilities of LLMs have been investigated for many different software engineering tasks, including library (i.e., non-web) API invocations (Zhuo et al., 2023), their aptitude for web API integrations appears to be rather unexplored (cf. also related work in Section 7). The question is: Given the pivotal role of web APIs, can we unlock the potential of LLMs for streamlining and accelerating the development processes of software that involves web API integration? As a first step towards answering this question, we need to understand how well existing LLMs support that task. To this end, we formulate the following research question:
RQ1: How correct is LLM-generated web API invocation code, and what types of errors are prevalent in such code?
Methodology for RQ1
@@
To arrive at an answer, we create WAPIIBench, a benchmark for web API integration tasks. Specifically, we perform the following steps. First, we create a dataset of web API invocation tasks and expected outcomes. Using a powerful state-of-the-art LLM, we generate the initial version of the dataset and subject it to a thorough manual review and correction process. Second, we implement an evaluation pipeline that allows us to safely execute web API invocation code and evaluate the resulting requests to obtain fine-grained correctness metrics. For this purpose, we develop a novel evaluation methodology. Third, we use our dataset and evaluation pipeline to evaluate a comprehensive set of LLMs, including models from the StarCoder (Li et al., 2023a; Lozhkov et al., 2024), DeepSeek-Coder (Guo et al., 2024b; DeepSeek-AI et al., 2024), Qwen-Coder (Hui et al., 2024), (Code) Llama (Rozière et al., 2023; Dubey et al., 2024), and GPT families.
@@
Listing 1: General structure of an API invocation using Axios and JavaScript
@@
axios.<method>(’https://server.com/path/to/endpoint/arg1’, {
@@
Since web APIs can be used in many ways, we need to be more specific and focus our research specifically on the following setup: We target real-world web APIs that adhere to OpenAPI [^1] https://www.openapis.org/, the most widely used specification standard for web APIs (SmartBear, 2020). We generate code in JavaScript, one of the most popular programming languages (Statista, 2024), and use Axios [^2] https://axios-http.com/, the leading JavaScript library for invoking web APIs (Greif and Burel, 2022), which allows explicit configuration of all components of an HTTP requests.
@@
The general structure of the API invocations we consider is shown in Listing 1. An invocation comprises an HTTP method (GET, POST, etc.), a URL with server address and (potentially parameterized) path to the endpoint, a request body, a request header, and query parameters. Such web API invocations have some unique challenges compared to function calls in a local codebase: (1) Operations are identified by the combination of a method and a long URL string, not just a simple function name; (2) there is not just one but multiple argument lists and the number of available arguments can be much higher while many arguments have complex, nested data types; and (3) web API endpoints are documented externally, limiting accessibility for LLMs, and the used specification format is much more complex than, e.g., that of documentation comments used for local functions.
Evaluation results for RQ1
@@
Our study reveals that while most LLMs demonstrate some understanding of the assigned tasks, show familiarity with the general structure of invocations in JavaScript/Axios, and appear to have memorized portions of the APIs’ specifications, they frequently generate incorrect invocations. In particular, they often hallucinate endpoint URLs and parameter names.
Based on our empirical findings, the central question shifts from whether LLMs can streamline and accelerate software development involving web API integration to how their performance on web API invocation tasks can be improved.
Two established techniques have been extensively investigated for improving the performance of LLMs in new domains: retrieval-augmented generation (RAG) (Lewis et al., 2020), which inserts relevant information into prompts, and fine-tuning, which adapts models through domain-specific training. While prior research (see Section 7) indicates both methods could prove effective in our setting, each approach has notable limitations. RAG risks prompt bloat (Gan and Sun, 2025), which consumes limited context window capacity, obscures critical information (Liu et al., 2024a), and increases computational costs. Fine-tuning requires substantial computational resources and risks obsolescence when API specifications evolve. Importantly, neither method guarantees correctness or code quality in generated outputs.
Given the limitations of RAG and fine-tuning approaches, this paper investigates constrained decoding (Hokamp and Liu, 2017; Deutsch et al., 2019) as an alternative method to improve reliability in LLM-generated web API integration code. Constrained decoding (also known as structured or guided generation) dynamically filters the token predictions of an LLM during generation to enforce compliance with predefined structural or semantic constraints. Crucially, this technique operates without requiring model modifications or prompt adjustments. To our knowledge, constrained decoding remains unexplored for web API invocations that explicitly configure all components of an HTTP request.
To systematically study the effectiveness of constrained decoding for improving the performance of LLMs on web API invocation tasks, we formulate and answer our second research question:
RQ2: How does constrained decoding influence the correctness of LLM-generated web API invocation code?
Methodology for RQ2
To answer RQ2, we perform two steps. First, we design a constrained decoding approach suitable for the specific use case of web API invocation tasks. Second, we use WAPIIBench to compare the performance of constrained and unconstrained LLMs.
The main challenge we have to address when applying constrained decoding in our setting is constraint creation. Since every web API requires different constraints, we implement an automatic translation from OpenAPI specifications to constraints that are represented as regular expressions. When generating the constraints, the special characteristics and complexities of web API invocations need to be taken into account. Besides respecting the general JavaScript and the specific Axios syntax (cf. Listing 1), our constraints need to handle interdependencies between allowed URLs, HTTP methods, and different kinds of parameters. Moreover, the number, ordering, and nesting of parameters can be arbitrary, while some parameters are required and some are optional, making advanced constraint formulations necessary.
Evaluation results for RQ2
The results of applying our evaluation methodology to study constrained models confirm that decoding with our constraints enforces compliance with an API’s specification when generating an invocation, reliably reducing the amount of illegal URLs, HTTP methods, and arguments to zero. Moreover, across models, constrained decoding significantly improves the overall correctness of the generated code, at relative gains of 90% and 142% on average depending on the provided starter code.
@@
In summary, by introducing WAPIIBench and exploring constrained decoding for web API invocations, we make the following contributions:
A first-of-its-kind dataset of web API invocation tasks:
across four real-world APIs with nearly 400 pairs of tasks and expected outcomes, which can be used for execution-based correctness analysis. → Section 2.1
An open-source evaluation pipeline:
for automatically and safely evaluating the performance of code generation models on web API invocation tasks using fine-grained metrics. → Section 2.2 – Section 2.4
A constraint generator:
that translates OpenAPI specifications to regex-based constraints, which ensure that generated code is specification-complaint. → Section 4
Novel empirical insights:
on the (in)ability of LLMs to generate correct API invocation code without and with constrained decoding. → Section 3 & Section 5
WAPIIBench is available on GitHub [^3] https://github.com/stg-tud/WAPIIBench; Note that the code also includes work-in-progress implementations for future work, which can be ignored.. In addition, we provide all model-generated codes in our artifact [^4] https://doi.org/10.5281/zenodo.13758414. Additional details and comprehensive results are provided in the appendix at the end of this paper.
2. Benchmark Design
Refer to caption Flow diagram visualizing the evaluation pipeline described in the caption.
Figure 1. Benchmark design for evaluating the capabilities of LLMs in generating web API invocation code. (1) Based on an API $a$ and its specifications $s$, an API invocation tasks $t$ and corresponding correct request configurations $c$ are created. (2) For each $t$, the LLM under evaluation generates an API invocation ii. (3) $i$ is executed in a controlled environment, yielding a request configuration $c^{\prime}$. (4) $c^{\prime}$ is compared to $c$ and validated against $s$ to obtain various metrics, shown in Table 2.
WAPIIBench’s evaluation pipeline, depicted in Figure 1, consists of four stages: (1) dataset creation, (2) code generation, (3) code execution, and (4) correctness analysis. They are detailed in the following subsections.
2.1. Dataset Creation
@@
Instead, we opted for a synthetic approach, using an LLM $\operatorname{Gen}$ and API specifications $s$ to generate task–configuration pairs for our dataset: $\operatorname{Gen}(a,s)=(t,c)$. To ensure our dataset represents real-world scenarios, we chose APIs of popular services across different domains—Asana, Google Calendar, Google Sheets, and Slack—which are specified following OpenAPI, the most commonly used industry standard (SmartBear, 2020). As $\operatorname{Gen}$, we used Gemini 1.5 Pro [^6] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/ for its very large context window (2M tokens), given that OpenAPI specifications can be very long (e.g., Slack’s contains over 20,000 lines of text). Together with the full specification, we provided the model with very detailed (but mostly non-API-specific) instructions on how to construct $t$ and $c$ in a zero-shot prompt to avoid pitfalls of few-shot prompting (Reynolds and McDonell, 2021) and increase generalizability. The prompt is provided in the appendix. In total, we synthesized 395 samples—one for each endpoint of the selected APIs. A detailed breakdown is shown in Table 1.
To ensure the validity of the AI-generated data, we curated all samples with automated and manual checks. The automated checks caught any inconsistencies between the generated configurations $c$ and the specifications $s$. The 9 samples that failed these checks were manually inspected and corrected. Afterward, we manually checked all 395 samples to catch remaining issues that cannot be checked automatically, especially underspecification and ambiguities in the tasks $t$, in order to ensure that there is one unique $c$ for each $t$. Concretely, we checked (1) that $t$ is well-formulated and solvable; (2) that $c$ actually solves $t$; and (3) that every argument in $t$ is reflected in $c$ and vice versa. 58 samples had one or multiple issues and were fixed by adjusting $t$ and/or $c$.
@@
Table 1. Number of endpoints per HTTP method for each API (each endpoint corresponds to one sample in the dataset)
API Total POST GET PUT DELETE PATCH
@@
Listing 2: Simplified task from our dataset with correct implementation
@@
axios.post(’https://www.googleapis.com/calendar/v3/calendars’, {
@@
Listing 3: Request configuration corresponding to the task in Listing 2
@@
"url": "https://www.googleapis.com/calendar/v3/calendars",
@@
2.2. Code Generation
Note that the code generation task described here is considerably different from the dataset generation task from Section 2.1. In the latter, Gemini had access to the full API specification and very detailed instructions, making this a task about processing in-context information. The models under evaluation, on the other hand, have to rely solely on memorized knowledge about the APIs, as we want to find out how well they can recall these details from their training data. This is a relevant concern as context length is limited, and it is often not desirable to fill the context with too much information. For instance, Slack’s specification has around 126K tokens, exceeding the context length of most models we evaluated.
2.3. Code Execution
Our approach to evaluating LLMs’ capabilities in generating API invocation code is based on functional correctness rather than syntactic similarity to a reference implementation, as the latter is prone to misjudgments (Evtikhiev et al., 2023). To measure functional correctness, we need to execute the generated code and assess its outcome, analogous to unit testing (Chen et al., 2021).
When executing $i$, we need to address challenges unique to our setting. First, direct execution of generated code is unsafe as it could send arbitrary requests to external servers. Second, we need to capture and evaluate the configuration $c^{\prime}$ of the API request implemented by $i$, which is a side effect and not the return value of $i$. Another common challenge when evaluating code generation is excess code (Guo et al., 2024a), i.e., code produced when the model terminates too late.
@@ 2.3. Code Execution
2.4. Correctness Analysis
Table 2. Evaluation Metrics for Correctness and Specification-Compliance of API Invocations
@@ 2.4. Correctness Analysis
We compare the configuration $c^{\prime}$, captured in the previous step, with the ground-truth $c$ from our dataset: $c^{\prime}\overset{?}{=}c$. However, as web API requests contain multiple components—URL, method, and parameters in different locations (cf. Listing 1)—only looking for exact matches is insufficient. Instead, we compare the configurations in an element-wise manner as well as in aggregated forms to facilitate a fine-grained evaluation. Additionally, we validate $c^{\prime}$ against the API specification $s$ to distinguish between endpoints and arguments that are undesired but specification-compliant and ones that are actually illegal. The key metrics we calculate are explained in Table 2 and additional ones are provided in the appendix. Note that some metrics apply only to one of the setups (full/argument completion), as mentioned in Section 2.2.
3. Evaluation of Unconstrained LLMs
Using our dataset and evaluation pipeline, we study the performance of 21 state-of-the-art open-source LLMs on web API invocation code generation. We primarily evaluated models that have undergone fine-tuning on code data—CodeT5+ (Wang et al., 2023) (16B), StarCoder (Li et al., 2023a) (15.5B), StarCoder2 (Lozhkov et al., 2024) (3B, 7B, 15B), DeepSeek-Coder (Guo et al., 2024b) (1.3B, 6.7B, 7B, 33B), DeepSeek-Coder-V2 (DeepSeek-AI et al., 2024) (16B), Qwen2.5-Coder (Hui et al., 2024) (0.5B, 1.5B, 3B, 7B, 14B, 32B), and Code Llama (Rozière et al., 2023) (7B, 13B, 70B)—and additionally Llama 3.1 (Dubey et al., 2024) (8B, 70B). These models were selected due to their frequent appearance in coding leaderboards [^7] E.g., https://evalplus.github.io/leaderboard.html. To obtain an upper bound on performance, we also evaluated OpenAI’s commercial GPT-4o [^8] https://openai.com/index/hello-gpt-4o/ and GPT-4o mini [^9] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ models. For a lack of space and for more visual clarity, we show only a subset of these models in our plots and tables, but our following analysis considers all of them. Comprehensive result tables are provided in the appendix.